OpenStudy (anonymous):

] Select the equations that are parallel and perpendicular to y = -3x - 1 and that pass through the point (3, 1).

OpenStudy (anonymous):

perpindicular=the slope is the negative reciprocal of the orginal line, i.e: 1/3x Parrallel=the slope is the same as the orginal line, i.e: -3x

OpenStudy (anonymous):

HERE ARE THE CHOICES: parallel: y = -3x + 6 perpendicular: y = x + parallel: y = -3x + 10 perpendicular: y = x parallel: y = x perpendicular: y = -3x parallel: y = x + 1 perpendicular: y = 3x - 1

OpenStudy (anonymous):

parallel: y = -3x + 6 perpendicular: y =1/3 x + 2 2/3 parallel: y = -3x + 10 perpendicular: y =1/3 x parallel: y =-1/3 x perpendicular: y = -3x parallel: y =1/3 x + 1 perpendicular: y = 3x - 1